# Accelerating the Search for Superconductors Using Machine Learning

### v1

@@ V.2 Using the compounds in Materials Project
To identify potential superconducting materials, we screen compounds from the Materials Project database [34], which hosts over 150,000 inorganic materials. We refine our selection by filtering for nonmagnetic, metallic, and experimentally reported compounds that are also present in the ICSD [50] and Pauling File database [73], as illustrated in Fig. 11. This reduces the dataset to approximately 15,352 compounds, of which around 1,606 are already listed in SuperCon-MTG, leaving approximately 13,746 candidate materials. We apply our machine learning model to this refined dataset to identify potential superconducting materials. The top 10 compounds with the highest predicted critical temperatures, as determined using both the 30-feature and 5-feature models, are presented in Table 5. 4 of these 10 compounds turn out to be reported superconductors, but not present in SuperCon-MTG. The predicted critical temperatures for $\text{Tl}_{2}\text{Ba}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$, $\text{Er}\text{Ba}_{2}\text{Cu}_{4}\text{O}_{8}$, and $\text{Tl}_{2}\text{Sr}_{2}\text{Ca}\text{Cu}_{2}\text{O}_{7}$ from both the 30-feature and 5-feature models are very close to the corresponding experimental values. For $\text{Bi}_{2}\text{Sr}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$, the prediction from the 30-feature model aligns well with the experimental $T_{c}$, while the 5-feature model underestimates it, though it still successfully identifies the compound as a superconductor. The remaining six materials are new predictions and should be tested for superconductivity. Most of the compounds listed in Table 5 contain copper in the +2 oxidation state. In $\text{Tl}_{5}\text{Ba}_{6}\text{Ca}_{6}\text{Cu}_{9}\text{O}_{29}$, copper exists in both +1 and +2 oxidation states. In $\text{Er}\text{Ba}_{2}\text{Cu}_{4}\text{O}_{8}$, the average oxidation state of copper is +2.25, with three copper atoms in the +2 state and one in the +3 state. The presence of copper in the +2 state suggests that hole doping is required for the material to exhibit superconductivity.
@@
Figure 11: A schematic of the filtering process applied to the Materials Project dataset in search of potential superconductors using machine learning models.
@@
mp-556574 $\text{Tl}_{2}\text{Ba}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$ 101.6 98.5 108 [74] Cuprates
@@
mp-1191679 $\text{Bi}_{2}\text{Sr}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$ 82.1 43.9 90 [75] Cuprates
@@
mp-6583 $\text{Er}\text{Ba}_{2}\text{Cu}_{4}\text{O}_{8}$ 74.5 71.5 78.5 [76] Cuprates
@@
mp-20824 $\text{Tl}\text{Sr}_{2}\text{Ca}\text{Cu}_{2}\text{O}_{7}$ 67 69 80 [77] Cuprates
@@ V.3 Unexplored Space of materials
Manual search for potential superconducting materials within the Materials Project database is rather a time intensive and impractical process. To streamline this process, we use the SMACT library [78] in Python to generate binary compositions consisting of one transition metal and one main group element (details provided in Supplementary Information, Sec. V [41]). Elements from the 7th period, noble gases, and radioactive elements such as Tc, Po, and At are excluded. The generated compositions are then validated based on fundamental chemical principles, ensuring charge neutrality and conditions on electronegativity.
Furthermore, we compute the probability of these generated compounds to exist using the given oxidation states of constituents (see Supplementary Information Sec VI [41]) and retain only those with a probability greater than zero. This filtering process reduces the initial 5,085 possible compositions to 292, of which 19 are already listed in SuperCon-MTG. We further refine the dataset of remaining 273 compounds by applying the filtering criteria outlined in Fig. 11. From this, we shortlist 11 materials, 4 of which have already been reported as superconductors, predominantly in studies published after 2016, as presented in Table 6. In addition, we propose 7 new compounds as potential superconducting materials, as shown in Table 7.
@@
657372 $\text{Ta}_{2}\text{Se}$ 4.2 8.1 3.8 [79] Others
646899 NiTe 2.7 2.6 1.5 [80]b Tellurides
238590 FeS 5.8 7 5 [81] Iron Chalcogenides
- NbS 4.4 6.7 5.6 [82] Others
@@ VI Summary
In this work, we present a workflow to modify and refine existing dataset of superconducting compounds and present a cleaner and curated version. Using ideas from Quantum Structure Diagrams, we demonstrate clustering of superconducting compounds by their classes in principal component plane revealing meaningful structure in the data. Building on this foundation, we developed a machine learning framework for efficient classification of materials as superconductor, and predict its $T_{c}$ directly from its chemical composition. Remarkably, the model requires neither crystallographic information nor physical theory-based features, relying solely on Quantum Structure Diagram based descriptors and Random forest classification and regression algorithms. Importantly, the model generalizes well beyond the SuperCon-MTG dataset, accurately predicting superconductivity in newly discovered compounds for superconductivity. We also identify several promising candidates for superconductivity, which will stimulate experiments to validate. This showcases our model’s potential to accelerate the discovery of novel superconducting materials.
@@ VII Acknowledgments
S.A. acknowledges JNCASR for a research fellowship. The authors thank Prof. Ram Seshadri for constructive suggestions and comments on manuscript. We thank Prof. A. Sundaresan and Prof. Ricardo Grau-Crespo for valuable discussions and insights.

### v2

@@ V.2 Using the compounds in the Materials Project
To identify potential superconducting materials, we screen compounds from the Materials Project database [Jain2013], which hosts over 150,000 inorganic materials. We refine our selection by filtering for nonmagnetic, metallic, and experimentally reported compounds that are also present in the ICSD [Belsky2002] and Pauling File database [Villars2019], as illustrated in Fig. 11. This reduces the dataset to approximately 15,352 compounds, of which around 1,606 are already listed in SuperCon-MTG, leaving approximately 13,746 candidate materials. We apply our machine learning model to this refined dataset to identify potential superconducting materials. The top 10 compounds with the highest predicted critical temperatures, as determined using both the 30-feature and 5-feature models, are presented in Table 5. 4 of these 10 compounds turn out to be reported superconductors, but not present in SuperCon-MTG. The predicted critical temperatures for $\text{Tl}_{2}\text{Ba}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$, $\text{Er}\text{Ba}_{2}\text{Cu}_{4}\text{O}_{8}$, and $\text{Tl}_{2}\text{Sr}_{2}\text{Ca}\text{Cu}_{2}\text{O}_{7}$ from both the 30-feature and 5-feature models are very close to the corresponding experimental values. For $\text{Bi}_{2}\text{Sr}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$, the prediction from the 30-feature model aligns well with the experimental $T_{c}$, while the 5-feature model underestimates it, though it still successfully identifies the compound as a superconductor. The remaining six materials are new predictions and should be tested for superconductivity. Most of the compounds listed in Table 5 contain copper in the +2 oxidation state. In $\text{Tl}_{5}\text{Ba}_{6}\text{Ca}_{6}\text{Cu}_{9}\text{O}_{29}$, copper exists in both +1 and +2 oxidation states. In $\text{Er}\text{Ba}_{2}\text{Cu}_{4}\text{O}_{8}$, the average oxidation state of copper is +2.25, with three copper atoms in the +2 state and one in the +3 state. The presence of copper in the +2 state suggests that hole doping is required for the material to exhibit superconductivity.
@@
Figure 11: A schematic of the filtering process applied to the Materials Project database in search of potential superconductors using machine learning models.
@@
mp-556574 $\text{Tl}_{2}\text{Ba}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$ 101.6 98.5 108 [Dingyun1989] Cuprates
@@
mp-1191679 $\text{Bi}_{2}\text{Sr}_{2}\text{Ca}_{3}\text{Cu}_{4}\text{O}_{12}$ 82.1 43.9 90 [Matheis1990] Cuprates
@@
mp-6583 $\text{Er}\text{Ba}_{2}\text{Cu}_{4}\text{O}_{8}$ 74.5 71.5 78.5 [Fischer1992] Cuprates
@@
mp-20824 $\text{Tl}\text{Sr}_{2}\text{Ca}\text{Cu}_{2}\text{O}_{7}$ 67 69 80 [AbdShukor1996] Cuprates
@@ V.3 Unexplored Space of materials
Manual search for potential superconducting materials within the Materials Project database is rather a time intensive and impractical process. To streamline this process, we use the SMACT library [Davies2019] in Python to generate binary compositions consisting of one transition metal and one main group element (details provided in Supplementary Information, Sec. VI [SI]). Elements from the 7th period, noble gases, and radioactive elements such as Tc, Po, and At are excluded. The generated compositions are then validated based on fundamental chemical principles, ensuring charge neutrality and conditions on electronegativity.
Furthermore, we compute the probability of these generated compounds to exist using the given oxidation states of constituents (see Supplementary Information Sec VII [SI]) and retain only those with a probability greater than zero. This filtering process helps to retain chemically plausible compositions and reduces the initial 5,085 possible compositions to 292, of which 19 are already listed in SuperCon-MTG. We further refine the dataset of remaining 273 compounds by applying the filtering criteria outlined in Fig. 11. From this, we shortlist 11 materials, 4 of which have already been reported as superconductors, predominantly in studies published after 2016, as presented in Table 6. In addition, we propose 7 new compounds as potential superconducting materials, as shown in Table 7.
@@
657372 $\text{Ta}_{2}\text{Se}$ 4.2 8.1 3.8 [Gui2020] Others
646899 NiTe 2.7 2.6 1.5 [Liang2023]b Tellurides
238590 FeS 5.8 7 5 [Lai2015] Iron Chalcogenides
- NbS 4.4 6.7 5.6 [Ruan2023] Others
@@ VI Summary
In this work, we present a workflow to modify and refine existing database of superconducting compounds and present a cleaner and curated version. Using ideas from Quantum Structure Diagrams, we demonstrate clustering of superconducting compounds by their classes in the plane of principal components revealing meaningful structure in the data. Building on this foundation, we developed a machine learning framework for efficient classification of materials as superconductor, and prediction of its $T_{c}$ directly from its chemical composition. Remarkably, the model requires no crystallographic information and relies only on Quantum Structure Diagram based descriptors and Random forest classification and regression algorithms. Importantly, the model generalizes well beyond the SuperCon-MTG dataset, accurately predicting superconductivity in newly discovered compounds for superconductivity. We also identify several promising candidates for superconductivity, which will stimulate new experiments. This showcases our model’s potential to accelerate the discovery of novel superconducting materials.
However, our model is limited by the lack of reliable non-superconductor data, which introduces dataset bias and constraints the development of more accurate $T_{c}$ prediction models. The use of composition-only descriptors further restricts the framework by omitting key physical factors underlying superconductivity. Incorporating crystal-structure information and synthesis conditions available from the literature, represents a natural next step toward building more accurate machine-learning models and advancing our understanding of superconductivity through materials informatics.
@@ VII Acknowledgments
S.A. acknowledges JNCASR for a research fellowship. The authors thank Prof. Ram Seshadri for constructive suggestions and comments on manuscript. We thank Prof. A. Sundaresan and Prof. Ricardo Grau-Crespo for valuable discussions and insights. U.V.W. acknowledges support from J. C. Bose Grant (2025/0210) of ANRF, Government of India.